A 20.0 mL sample of 0.150M ethylamine is titrated with 0.0981 M HCl. What is the pH after the addition of 5.0 mL of HCI? For ethylamine, pkb =3.25.
Answer #1

Initial moles of C2H5NH2 = 20/1000 x 0.150 = 0.003 mol Moles of HCl added = 5/1000 x 0.0981 = 0.0004905 mol Moles of C2H5NH2